# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ III-A5 Finance
For higher effectiveness and efficiency, a transformer model can be asymmetrically constructed with a shallower encoder and a deeper decoder. To achieve better performances, it is necessary to employ strategies such as massively scaling up sampling, followed by the filtering and clustering of samples into a compact set. The utilization of novel sampling-efficient transformer architectures designed to facilitate large-scale sampling is crucial. Simplifying problem descriptions can effectively improve the model’s performance.
@@
Models Findings & Insights
Chinchilla
The experiments that culminated in the development of Chinchilla determined that for optimal computation during training, the model size and the number of training tokens should be scaled proportionately: for each doubling of the model size, the number of training tokens should be doubled as well.
PaLM
English-centric models produce better translations when translating to English as compared to non-English Generalized models can have equivalent performance for language translation to specialized small models Larger models have a higher percentage of training data memorization Performance has not yet saturated even at 540B scale, which means larger models are likely to perform better
AlexaTM
Compared to commonly used Decoder-only Transformer models, seq2seq architecture is more suitable for training generative LLMs given stronger bidirectional attention to the context. An extra Causal Language Modeling (CLM) task can be added to benefit the model with a more efficient in-context learning, especially for few-shot learning tasks. The key to training powerful seq2seq-based LLMs lies in mixed pre-training, rather than additional multitask training. Placing layernorms at the beginning of each transformer layer can improve the training stability of large models.
@@
Training with a mixture of denoisers outperforms PaLM when trained further for a few more FLOPs Training with a mixture of denoisers improves the infilling ability and open-ended text generation diversity
@@
Mode switching training enables better performance on downstream tasks CoT prompting outperforms standard prompting for UL2
@@
Pre-training data with a small proportion of multi-task instruction data improves the overall model performance
@@
Multi-step prompting for code synthesis leads to a better user intent understanding and code generation
@@
LLaMA is open-source and can be fine-tuned or continually pre-trained to develop new models or instruction-based tools. A few optimizations are proposed to improve the training efficiency of LLaMA, such as efficient implementation of multi-head self-attention and a reduced amount of activations during back-propagation. Training exclusively on public data can also achieve state-of-the-art performance. A constant performance improvement is gained when scaling the model. Smaller models can also realize good performances using more training data and time.
@@
Sparse models provide the benefits of large models at a lower computation cost Randomly Routed Experts reduces catastrophic forgetting effects which in turn is essential for continual learning Randomly Routed Experts allow extracting a domain-specific sub-model in deployment which is cost-efficient while maintaining a performance similar to the original
@@
Pre-training with general-purpose and task-specific data improves task performance without hurting other model capabilities
@@
Combining pre-training and fine-tuning stages in single training avoids catastrophic forgetting
@@
Causal LM is crucial for a model’s generation capability in encoder-decoder architectures Multiple training objectives like span corruption, Causal LM, matching, etc complement each other for better performance
@@
HHH prompt by Anthropic allows the model to follow instructions without fine-tuning
@@
Model trained on unfiltered data is more toxic but may perform better on downstream tasks after fine-tuning Model trained on unfiltered data requires fewer samples for safety alignment
@@
Data quality is important to train better models Model and data size should be scaled with 1:1 proportions Smaller models trained for larger iterations outperform larger models
TABLE II: Key insights and findings from the study of instruction-tuned Large Language Models.
Models Findings & Insights
@@
Multi-task prompting enables zero-shot generalization and outperforms baselines Even a single prompt per dataset task is enough to improve performance
@@
The answer quality of LLMs can be further improved with human feedback. To aid the model in effectively filtering and utilizing relevant information, human labelers play a crucial role in answering questions regarding the usefulness of the retrieved documents. Interacting a fine-tuned language model with a text-based web-browsing environment can improve end-to-end retrieval and synthesis via imitation learning and reinforcement learning. Generating answers with references can make labelers easily judge the factual accuracy of answers.
@@
Instruction tuning leads to a stronger generalization of unseen tasks More tasks improve generalization whereas only increasing task instances does not help Supervised trained models are better than generalized models Models pre-trained with instructions and examples perform well for different types of inputs
@@
Instruction tuning enables zero-shot generalization to the tasks never seen before Multi-lingual training leads to even better zero-shot generalization for both English and non-English Training on machine-translated prompts improves performance for held-out tasks with non-English prompts English only fine-tuning on multilingual pre-trained language model is enough to generalize to other pre-trained language tasks
@@
Task size sampling to create a batch with most of the task examples is important for better performance Only example proportional sampling is not enough, training datasets/benchmarks should also be proportional for better generalization/performance Fully held-out and partially supervised tasks performance improves by scaling tasks or categories whereas fully supervised tasks have no effect Including small amounts i.e. 5% of pretraining data during fine-tuning is effective Only 1% reasoning data improves the performance, adding more deteriorates performance Adding dialogue data makes the performance worse
@@
Finetuning with CoT improves performance on held-out tasks Fine-tuning along with CoT data improves reasoning abilities CoT tuning improves zero-shot reasoning Performance improves with more tasks Instruction fine-tuning improves usability which otherwise is challenging for pre-trained models Improving the model’s performance with instruction tuning is compute-efficient Multitask prompting enables zero-shot generalization abilities in LLM
Sparrow
The judgments of labelers and the alignments with defined rules can help the model generate better responses. Good dialogue goals can be broken down into detailed natural language rules for the agent and the raters. The combination of reinforcement learning (RL) with reranking yields optimal performance in terms of preference win rates and resilience against adversarial probing.
@@
Fine-tuning with re-written instruction-tuning data into a complex set improves the performance significantly
@@
Model learns to write safe responses with fine-tuning on safe demonstrations, while additional RLHF step further improves model safety and make it less prone to jailbreak attacks
@@
Less high quality data is enough for fine-tuned model generalization
III-B Fine-Tuned LLMs
@@
III-B1 Instruction-Tuning with Manually Created Datasets
@@
III-B2 Instruction-Tuning with LLMs Generated Datasets
@@
LLaMA Tuned: Various models in literature instruction-tune LLaMA [145] with GPT-3 [6] or GPT-4 [146] generated datasets. Among these, Alpaca [147], Vicuna [148], and LLaMA-GPT-4 [149] are a few general-purpose fine-tuned models, where Alpaca is trained on 52k samples from text-davinci-003, Vicuna on 70k samples from ShareGPT.com, and LLaMA-GPT-4 by re-creating Alpaca instructions from GPT-4. Goat [150] fine-tunes LLaMA for arithmetic tasks (1 million samples) by generating data from ChatGPT and outperforms GPT-4, PaLM, BLOOM, OPT, etc, attributing its success to the LLaMA’s consistent tokenization of numbers. HuaTuo [151] is a medical knowledge model, fine-tuned with a generated QA dataset of 8k instructions.
@@
III-B3 Aligning with Human Preferences
@@
  Aligning with Prompts: LLMs can be steered with prompts to generate desirable responses without training [164, 165]. The self-correction prompting in [165] concatenates instructions and CoT with questions, guiding the model to answer its instruction following strategy to ensure moral safety before the actual answer. This strategy is shown to reduce the harm in generated responses significantly.
  Red-Teaming/Jailbreaking/Adversarial Attacks: LLMs exhibit harmful behaviors, hallucinations, leaking personal information, and other shortcomings through adversarial probing. The models are susceptible to generating harmful responses even though they are aligned for safety [166, 167]. Red-teaming is a common approach to address illicit outputs, where the LLMs are prompted to generate harmful outputs [167, 168]. The dataset collected through red-teaming is used to fine-tune models for safety. While red-teaming largely relies on human annotators, another work [169] red-team LLMs to find prompts that lead to harmful outputs of other LLMs.
III-B4 Continue Pre-Training
@@
III-B5 Sample Efficiency
While fine-tuning data is generally many-fold smaller than the pre-training data, it still has to be large enough for acceptable performance [16, 92, 18] and requires proportional computing resources. To study the effects on performance with less data, existing literature [172, 173] finds that the models trained on lesser data can outperform models trained with more data. In [172], 25% of the total downstream data is found enough for state-of-the-art performance. Selecting coreset-based 0.5% of the total instruction-tuning data improves the model performance by 2% in [173], as compared to the complete data tuning. Less is more for alignment (LIMA) [174] uses only 1000 carefully created demonstrations to fine-tune the model and has achieved comparable performance to GPT-4.
III-C Increasing Context Window
@@
Position Interpolation: Rather than extrapolating, [49] shows that interpolating position encodings within the pre-trained context window are more effective. The work demonstrates that only 1000 steps of fine-tuning are enough to achieve better results on larger windows without performance loss compared to the original context size. Giraffe [46] uses power scaling in RoPE, and YaRN [47] proposed NTK-aware interpolation.
Efficient Attention Mechanism: Dense global attention is one of the major constraints in training larger context window LLMs. Using efficient attention variants, such as local, sparse, and dilated attention, reduces the computation cost significantly. LongT5 [48] proposes transient global attention (TGlobal), applying attention to local and global tokens (windowing token averaging). The model replaces attention in T5 [10] with TGlobal attention, pre-trains the model on 4098 sequence length, fine-tunes on larger window sizes, as large as 16k, and improves task performance with longer inputs. This shows the extrapolation ability of TGlobal attention with only fine-tuning. COLT5 [176] uses two branches, one with lightweight and the other with heavyweight attention and feed-forward layers. All tokens are processed from the lightweight branch, and only important tokens are routed to the heavyweight branch. LongNet [177] replaces standard attention with dilated attention, expanding sequence length to 1 billion tokens. LongLoRA [178] proposes shift-short attention, used during fine-tuning to reduce dense attention costs, while the model during inference can use dense attention and achieve similar performance as full attention fine-tuning.
@@
III-D Augmented LLMs
LLMs are capable of learning from the examples concatenated with the input, known as context augmentation, in-context learning (ICL), or few-shot prompting. They show excellent generalization to unseen tasks with few-shot prompting, enabling LLMs to answer queries beyond the capacity acquired during training [6, 55]. These emergent abilities allow for adapting the model without fine-tuning - a costly process. Aside from this, hallucination, producing inaccurate, unsafe, or factually incorrect responses, is common for LLMs, which is avoided by augmenting contextual data. While the user can provide in-context samples in the query [54, 32], here we specifically refer to the methods that access external storage programmatically, calling them augmented LLMs.
@@ III-D Augmented LLMs
III-D1 Retrieval Augmented LLMs
@@
III-D2 Tool Augmented LLMs
@@
Training with Tool Augmentation: LLMs are trained to interact with diverse tools, enhancing planning abilities to overcome the limitations of zero-shot tool augmentation [210, 27, 211, 212]. Gorilla [210] instruction-tunes LLaMA with information retrieval from API documentation. It uses self-instruct [19] data generation pipeline with GPT-4 by providing in-context examples retrieved from API documentation. Tool augmented language model (TALM) [27] fine-tunes T5 [10] for tool use with a self-play approach, where it iteratively completes tool manipulation tasks and includes them back in the training set. ToolLLM [212] collects 16k APIs from RapidAPI. It samples APIs from the list to generate an instruction-tuning dataset using ChatGPT in single-tool and multi-tool scenarios. For high-quality datasets, ToolLLM suggested a depth-first search-based decision tree (DFSDT) method to generate ground-truths with diverse reasoning and planning.
[…]
III-E LLMs-Powered Agents

### 09 Apr 2024 v9

@@ 1. The model can be fine-tuned to learn to call different external information resources and tools
1. For higher effectiveness and efficiency, a transformer model can be asymmetrically constructed with a shallower encoder and a deeper decoder 2. To achieve better performances, it is necessary to employ strategies such as massively scaling upsampling, followed by the filtering and clustering of samples into a compact set 3. The utilization of novel sampling-efficient transformer architectures designed to facilitate large-scale sampling is crucial 4. Simplifying problem descriptions can effectively improve the model’s performance
Chinchilla
1. The model size and the number of training tokens should be scaled proportionately: for each doubling of the model size, the number of training tokens should be doubled as well
PaLM
1. English-centric models produce better translations when translating to English as compared to non-English 2. Generalized models can have equivalent performance for language translation to specialized small models 3. Larger models have a higher percentage of training data memorization 4. Performance has not yet saturated even at 540B scale, which means larger models are likely to perform better
AlexaTM
1. Encoder-decoder architecture is more suitable to train LLMs given bidirectional attention to the context than decoder-only 2. Causal Language Modeling (CLM) task can be added to benefit the model with efficient in-context learning 3. Placing layer norm at the beginning of each transformer layer improves the training stability
@@
[…]
@@
1. Training with a mixture of denoisers outperforms PaLM when trained further for a few more FLOPs 2. Training with a mixture of denoisers improves the infilling ability and open-ended text generation diversity
@@
1. Mode switching training enables better performance on downstream tasks 2. CoT prompting outperforms standard prompting for UL2
@@
1. Pre-training data with a small proportion of multi-task instruction data improves the overall model performance
@@
1. Multi-step prompting for code synthesis leads to a better user intent understanding and code generation
@@
1. A constant performance improvement is observed when scaling the model 2. Smaller models can achieve good performances with more training data and computing time
@@
1. Sparse models provide the benefits of large models at a lower computation cost 2. Randomly Routed Experts reduces catastrophic forgetting effects which in turn is essential for continual learning 3. Randomly Routed Experts allow extracting a domain-specific sub-model in deployment which is cost-efficient while maintaining a performance similar to the original
@@
1. Pre-training with general-purpose and task-specific data improves task performance without hurting other model capabilities
@@
1. Combining pre-training and fine-tuning stages in single training avoids catastrophic forgetting
@@
1. Causal LM is crucial for a model’s generation capability in encoder-decoder architectures 2. Multiple training objectives like span corruption, Causal LM, matching, etc complement each other for better performance
@@
1. HHH prompt by Anthropic allows the model to follow instructions without fine-tuning
@@
1. Model trained on unfiltered data is more toxic but may perform better on downstream tasks after fine-tuning 2. Model trained on unfiltered data requires fewer samples for safety alignment
@@
1. Data quality is important to train better models 2. Model and data size should be scaled with 1:1 proportions 3. Smaller models trained for larger iterations outperform larger models
Table 2: Key insights and findings from the study of instruction-tuned Large Language Models.
[…]
@@
1. Multi-task prompting enables zero-shot generalization and outperforms baselines 2. Even a single prompt per dataset task is enough to improve performance
@@
1. To aid the model in effectively filtering and utilizing relevant information, human labelers play a crucial role in answering questions regarding the usefulness of the retrieved documents 2. Interacting a fine-tuned language model with a text-based web-browsing environment can improve end-to-end retrieval and synthesis via imitation learning and reinforcement learning 3. Generating answers with references can make labelers easily judge the factual accuracy of answers
@@
1. Instruction tuning leads to a stronger generalization of unseen tasks 2. More tasks improve generalization whereas only increasing task instances does not help 3. Supervised trained models are better than generalized models 4. Models pre-trained with instructions and examples perform well for different types of inputs
@@
1. Instruction tuning enables zero-shot generalization to tasks never seen before 2. Multi-lingual training leads to even better zero-shot generalization for both English and non-English 3. Training on machine-translated prompts improves performance for held-out tasks with non-English prompts 4. English only fine-tuning on multilingual pre-trained language model is enough to generalize to other pre-trained language tasks
@@
1. Creating a batch with multiple task examples is important for better performance 2. Only example proportional sampling is not enough, training datasets should also be proportional for better generalization/performance 3. Fully held-out and partially supervised tasks performance improves by scaling tasks or categories whereas fully supervised tasks have no effect 4. Including small amounts i.e. 5% of pretraining data during fine-tuning is effective 5. Only 1% reasoning data improves the performance, adding more deteriorates performance 6. Adding dialogue data makes the performance worse
Sparrow
1. Labelers’ judgment and well-defined alignment rules help the model generate better responses 2. Good dialogue goals can be broken down into detailed natural language rules for the agent and the raters 3. The combination of reinforcement learning (RL) with reranking yields optimal performance in terms of preference win rates and resilience against adversarial probing
@@
1. Finetuning with CoT improves performance on held-out tasks 2. Fine-tuning along with CoT data improves reasoning abilities 3. CoT tuning improves zero-shot reasoning 4. Performance improves with more tasks 5. Instruction fine-tuning improves usability which otherwise is challenging for pre-trained models 6. Improving the model’s performance with instruction tuning is compute-efficient 7. Multitask prompting enables zero-shot generalization abilities in LLM
@@
1. Fine-tuning with re-written instruction-tuning data into a complex set improves performance
@@
1. Model learns to write safe responses with fine-tuning on safe demonstrations, while additional RLHF step further improves model safety and make it less prone to jailbreak attacks
@@
1. Less high quality data is enough for fine-tuned model generalization
3.2 Fine-Tuned LLMs
@@
3.2.1 Instruction-Tuning with Manually Created Datasets
@@
3.2.2 Instruction-Tuning with LLMs Generated Datasets
@@
LLaMA Tuned: Various models in the literature instruction-tune LLaMA [146] with GPT-3 [6] or GPT-4 [147] generated datasets. Among these, Alpaca [148], Vicuna [149], and LLaMA-GPT-4 [150] are a few general-purpose fine-tuned models, where Alpaca is trained on 52k samples from text-davinci-003, Vicuna on 70k samples from ShareGPT.com, and LLaMA-GPT-4 by re-creating Alpaca instructions from GPT-4. Goat [151] fine-tunes LLaMA for arithmetic tasks (1 million samples) by generating data from ChatGPT and outperforms GPT-4, PaLM, BLOOM, OPT, etc., attributing its success to the LLaMA’s consistent tokenization of numbers. HuaTuo [152] is a medical knowledge model, fine-tuned with a generated QA dataset of 8k instructions.
@@
3.2.3 Aligning with Human Preferences
@@
  Aligning with Prompts: LLMs can be steered with prompts to generate desirable responses without training [165, 166]. The self-correction prompting in [166] concatenates instructions and CoT with questions, guiding the model to answer its instruction following a strategy to ensure moral safety before the actual answer. This strategy is shown to reduce the harm in generated responses significantly.
  Red-Teaming/Jailbreaking/Adversarial Attacks: LLMs exhibit harmful behaviors, hallucinations, leaking personal information, and other shortcomings through adversarial probing. The models are susceptible to generating harmful responses even though they are aligned for safety [167, 168]. Red-teaming is a common approach to address illicit outputs, where the LLMs are prompted to generate harmful outputs [168, 169]. The dataset collected through red-teaming is used to fine-tune models for safety. While red-teaming largely relies on human annotators, another work [170] red-team LLMs to find prompts that lead to harmful outputs for other LLMs.
3.2.4 Continue Pre-Training
@@
3.2.5 Sample Efficiency
While fine-tuning data is generally many-fold smaller than the pre-training data, it still has to be large enough for acceptable performance [16, 97, 18] and requires proportional computing resources. Studying the effects on performance with less data, existing literature [173, 174] finds that models trained on less data can outperform models trained with more data. In [173], 25% of the total downstream data is found enough for state-of-the-art performance. Selecting coreset-based 0.5% of the total instruction-tuning data improves the model performance by 2% in [174], as compared to the complete data tuning. Less is more for alignment (LIMA) [175] uses only 1000 carefully created demonstrations to fine-tune the model and has achieved comparable performance to GPT-4.
3.3 Increasing Context Window
@@
Position Interpolation: Rather than extrapolating, [49] shows that interpolating position encodings within the pre-trained context window are more effective. The work demonstrates that only 1000 steps of fine-tuning are enough to achieve better results on larger windows without reducing performance compared to the original context size. Giraffe [46] uses power scaling in RoPE, and YaRN [47] proposed NTK-aware interpolation.
Efficient Attention Mechanism: Dense global attention is one of the major constraints in training larger context window LLMs. Using efficient attention variants, such as local, sparse, and dilated attention, reduces the computation cost significantly. LongT5 [48] proposes transient global attention (TGlobal), applying attention to local and global tokens (windowed token averaging). The model replaces attention in T5 [10] with TGlobal attention, pre-trains the model on 4098 sequence length, fine-tunes on larger window sizes, as large as 16k, and improves task performance on longer inputs. This shows the extrapolation ability of TGlobal attention with only fine-tuning. COLT5 [177] uses two branches, one with lightweight and the other with heavyweight attention and feed-forward layers. All tokens are processed from the lightweight branch, and only important tokens are routed to the heavyweight branch. LongNet [178] replaces standard attention with dilated attention, expanding sequence length to 1 billion tokens. LongLoRA [179] proposes shift-short attention, used during fine-tuning to reduce dense attention costs. However, the model during inference uses dense attention and achieves similar performance as full attention fine-tuning.
@@
3.4 Augmented LLMs
LLMs are capable of learning from the examples concatenated with the input, known as context augmentation, in-context learning (ICL), or few-shot prompting. They show excellent generalization to unseen tasks with few-shot prompting, enabling LLMs to answer queries beyond the capacity acquired during training [6, 55]. These emergent abilities allow for adapting the model without fine-tuning—a costly process. Aside from this, hallucination, producing inaccurate, unsafe, or factually incorrect responses, is common for LLMs, which is avoided by augmenting contextual data. While the user can provide in-context samples in the query [54, 32], here we specifically refer to the methods that access external storage programmatically, calling them augmented LLMs.
@@ 3.4 Augmented LLMs
3.4.1 Retrieval Augmented LLMs
@@
3.4.2 Tool Augmented LLMs
@@
Training with Tool Augmentation: LLMs are trained to interact with diverse tools, enhancing planning abilities to overcome the limitations of zero-shot tool augmentation [211, 27, 212, 213]. Gorilla [211] instruction-tunes LLaMA with information retrieval from API documentation. It uses the self-instruct [19] data generation pipeline with GPT-4 by providing in-context examples retrieved from API documentation. Tool augmented language model (TALM) [27] fine-tunes T5 [10] for tool use with a self-play approach, where it iteratively completes tool manipulation tasks and includes them back in the training set. ToolLLM [213] collects 16k APIs from RapidAPI. It samples APIs from the list to generate an instruction-tuning dataset using ChatGPT in single-tool and multi-tool scenarios. For high-quality datasets, ToolLLM suggested a depth-first search-based decision tree (DFSDT) method to generate ground-truths with diverse reasoning and planning.
Multimodal Tool Augmentation: The compositional reasoning capacity of LLMs allows them to manipulate tools in multimodal settings [205, 206, 214]. Following the pipeline shown in Figure 13, the LLM outlines a plan, generally executing in a sequence: Plan $\to\,$ Tool selection $\to\,$ Execute $\to\,$ Inspect $\to\,$ Generate, to respond to the user query. Here, the database of tools is rich in modalities, including text, images, etc. Many of the multimodal tool augmentation systems employ multimodal LLMs [31, 215, 214, 206], while others utilize single modality LLMs and generate a plan on using different modality tools to solve multimodal queries [216].
3.5 LLMs-Powered Agents